I am questioning if there is enough information to solve this problem:How much water should be added to 77 units of a 37% salt solution to get a 30% salt solution?
Thank you
Janet

Answer by stanbon(75887)

How much water should be added to 77 units of a 37% salt solution to get a 30% salt solution?
-----
Equation:
water + water = water
0.63*77 + x = 0.70(77+x)
---
Multiply thru by 100 to get:
63*77 + 100x = 70*77 + 70x
---
30x = 7*77
---
x = 17.97 units (# of units of water that are needed for the mixture)
==========================
You could also do the problem using the following idea:
salt + salt = salt
0.37*77 + 0*x = 0.30(77+x)
37*77 + 0 = 30*77+30x
7*77 = 30x
x = 7.77
==========================
Cheers,
Stan H.
====================